OpenStudy (anonymous):

How many and of what type are the solutions to x^2 + 7x + 10 = 0?

OpenStudy (anonymous):

well of course since the degree i.e. the highest power of the equation is 2, i.e. in x^2 , it has 2 solutions. Now do you know how to find b^2-4ac?

OpenStudy (anonymous):

ummm no @luckythebest

OpenStudy (anonymous):

Ok there is a small concept then. So you know that any quadratic equation can be written in the form of ax^2+bx+c = 0 ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

yep. so if 1x^2 + 7x + 10 = ax^2 + bx + c, a = 1, b = 7, c = 10. It is observed that finding b^2-4ac can help us to find the nature of the solutions of this equation. So if you can find b^2-4ac ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yep so how much does it come to

OpenStudy (anonymous):

-5, -2

OpenStudy (anonymous):

and 0

OpenStudy (anonymous):

lol those are the solutions. :P

OpenStudy (anonymous):

anyway; -5, -2 are unequal, rational, and real. = nature of solution

OpenStudy (anonymous):

so its Two different rational solutions?

OpenStudy (anonymous):

@luckythebest

OpenStudy (anonymous):

@allie_bear22 you got that right.